RC meeting and Run Plan, Friday March 10th
Return to the RC Meeting Calendar
Contents
RC Report
- RC: Holly
Summary
- RC Daily report: https://logbooks.jlab.org/entry/4146767
We took production for most of shifts, requesting and getting 45uA due to Hall C being down due to magnet troubles. We did suffer several DAQ issues throughout the past day which Sean summarized here: https://logbooks.jlab.org/entry/4146751. We took the SBS GEMs out of the DAQ for a few hours after their issues were unresolved by rebooting the crates, and they seemed to start magically working when we put them back in the DAQ earlier this morning. The beam went down at 10am this morning to do some RF repair work on a few modules in the south linac so that we can run with a better margin throughout the weekend. This work is estimated to be complete in 2-4 hours.
Swing summary: https://logbooks.jlab.org/entry/4146158 Owl summary: https://logbooks.jlab.org/entry/4146472
Recent Issues
- Golden run was updated last night using 30uA data https://logbooks.jlab.org/entry/4146241
Current Status
- Latest carbon hole image: https://logbooks.jlab.org/entry/4145626
- Running 45uA on pol-3He
Run Plan
- Production on 3He!
- GENII-3Stream config
- PS1=2, PS3=0, PS5=6
- Useful guide from Bill on rates to compare when bringing beam back: https://logbooks.jlab.org/entry/4144886
Data Collection and Polarization Status
- Currently 40%: https://logbooks.jlab.org/entry/4146672
- Data status at 4th-pass: 47 C (uncorrected)